Group Selection Page HELP Controls Group ITDM Vacuum WP78 SPB SQS XO ToBeMigrated2XWiki ToBeMigrated2GitlabPages ToBeMigratedByGroup Migrated2Zulip Archived2PDF not2MigrateOrArchive
SCS SXP Migrated2Zulip Photon Commissioning Team XRO DOC HED Migrated2Zulip WP74 MID PSPO_XO XO_Operation SQS Migrated2Zulip FXE Migrated2Zulip DET Sample Environment SPB Migrated2Zulip
OPERATION-2025 OPERATION-2024 OPERATION-2023 OPERATION-2022 OPERATION-2021 OPERATION-2020 OPERATION-2019 SASE1-2018 SASE1-2017 Experiment data
  OPERATION-2023 logbook, Contact: F.Wolff-Fabris, A.Galler, L.Samoylova, H.Sinn  Not logged in ELOG logo
Message ID: 219     Entry time: 02 Jun 2023, 23:34
Author: Romain Letrun 
Group: SPB 
Subject: Shift summary 

X-ray delivery

  • 10.65 keV, 128 pulses at 1.13 MHz, 1.6-1.7 mJ.
  • Overall stable delivery.

Optical laser delivery

  • 800 nm, 50 fs, 226 pulses at 1.13 MHz.

Achievements/Observations

  • Continued characterisation of crystals (thermal measurement, interferometry, imaging) under different beam conditions ranging from direct to focused beam.
  • Spectral measurement in parallel to crystals characterisation.
  • Imaging of static samples with Zyla and Shimadzu.

Issues

  • We were not able to complete the last part of our programme since we were not allowed to change the number of bunches and repetition rate. Measurement ended at 21:00 and last two hours fo the shift were spent in discussion with SQS/BKR/RC/PRC with no positive outcome.
ELOG V3.1.4-7c3fd00